# compat=1.1.0 New lower background and higher rate technique for anti-neutrino detection using Tungsten 183 Isotope

### v2

@@ Abstract
Low energy anti-neutrinos detected from reactors or other sources have typically used the conversion of an anti-neutrino on Hydrogen, producing a positron and a free neutron. This neutron is subsequently captured on a secondary element with a large neutron capture cross section such as Gadolinium or Cadmium. We have studied the anti-neutrino conversion and suggest other elements that have a comparable cross section for anti-neutrino reactions. With most neutron captures on Gadolinium, it is possible to get two or three delayed gamma signals of known energy to occur. Experiments like ATLAS can make measurements with timing on the order of 25 ns. With electronics like these, this leads to the possibility of having a triple delayed coincidence using the positron annihilation on atomic shell electrons as the starting signal. We have also found an isotope of Tungsten, W that offers a large anti-neutrino interaction cross section of m and an anti-neutrino threshold energy for the production of the ground state at 2.094 MeV and the production of the first excited state at 2.167 MeV. This reaction makes a nuclear m1 excited state of Ta that emits a signature secondary gamma pulse of 73 keV with a 106 ns half-life. This offers a new delayed coincidence technique that can be used to identify anti-neutrinos with lower background noise. This allows for less shielding than required for modern inverse beta decay detectors.
@@ 1 Introduction
Historically the first direct detection of the neutrino was using anti-neutrinos from a nuclear reactor by Cowan and Reines in the 1950s [1]. Nuclear reactors produce many different energies of anti-neutrinos, not just from the fission process itself, but also from the daughter nuclei produced in the fission, see Figure 1 [2]. Recently in cooperation with NASA funding is an idea to make a spacecraft with a solar neutrino detector [3]. That work uses a double pulse signal from a Gallium interaction, which was seminal in the studies of this paper. Instead of using double pulses from solar neutrinos, we have studied elements which could produce multiple coincidence from reactor anti-neutrinos, which led us to Tungsten 183.
The standard process of anti-neutrino detection[1], depicted in Figure 2 is a weak-nuclear force process of an anti-neutrino interacting with a proton:
where the positron annihilates on an atomic shell electron making a distinct signal via the positron annihilation, but the neutron is fast and must be slowed. The neutron’s passage through matter slows it to capturable energies, but it takes over 200 s. When the neutron is absorbed on a heavy element such as Gadolinium or Cadmium, the neutron emits many different possible cases of secondary gammas, see Table 1 for Gadolinium. One such process only involves the emission of a single gamma and the other 11 have two or more gammas emitted with delayed timings between them. This double coincidence was measured in the 1950s with slow old electronics of that era ( s [1]), but the large >200 s signal delay made that measurement possible. Today with fast electronics it is possible to consider using a triple delayed coincidence with smart electronics selection and reject out of time background signals. These electronics can be combined with fast scintillators like NaPGaW, with a decay timing between 20 and 32 s [4], or PbWO, with a decay timing < 10 ns [5]. If the anti-neutrino has enough energy this reaction can occur on a proton in a heavy element. With sufficient energy, the interaction can knock out a free neutron which, if the emitted neutron is low enough energy, can slow faster than the Hydrogen reaction giving shorter coincidence timing.
A key idea studied here was a comprehensive survey of the known atomic isotopes to find the most promising candidates. In section 2, we describe the search for elemental isotopes which have a large anti-neutrino cross section and our determination of their capacity to produce a delayed coincidence signal topology. We calculate anti-neutrino energy threshold and secondary emission parameters for a particularly promising process, which is described in section 3. Values chosen in this section are based on the typical power generated by small nuclear reactors, such as on nuclear submarines. We conclude in section 4 with a discussion of possible applications for this technology, including scientific, economic, and national security.
Primary | Secondary | Tertiary | Intensity [% 10] |
@@
2 Estimating Low Energy Anti-neutrino Interaction Cross Sections for The Most Promising Anti-neutrino Targets
For our work, we needed to determine which anti-neutrino targets are the most promising for detection in the low-energy regimes of reactor anti-neutrinos. Codes like GENIE [6] are accurate as low as 10 MeV or 50 MeV, depending upon the anti-neutrino target, but reactor anti-neutrinos using inverse beta decay with a free neutron have gone as low as 1.75 MeV. A good anti-neutrino target would be one which has the largest possible anti-neutrino cross section and the lowest possible anti-neutrino energy threshold. Using Genie, we have studied the available data and we have listed the isotopes with the ten highest cross sections for anti-neutrino interactions and their cross sections in Table 2 [7]. All of these cross sections are taken directly from the lowest energy point of the indicated GENIE physics model. We found that certain isotopes of Gold and Tungsten have a favorable cross-section, and one isotope of Tungsten can produce a fast double delayed coincidence, which will be described with more detail in section 3.
Model | Energy (MeV) | (b) | () | Isotope |
5290 | 10 | H | ||
4034 | 10 | Pb | ||
4029 | 10 | Au | ||
4024 | 10 | W | ||
4019 | 10 | W | ||
4004 | 10 | Xe | ||
3999 | 10 | Sn | ||
13342 | 10 | H | ||
13452 | 55.33 | Gd | ||
13448 | 52.83 | Ba |
If these especially high anti-neutrino cross section anti-neutrino targets can be incorporated into a fast scintillating crystal or glass then they can be made into a detector system with applications for science measurements, imaging and surveillance apparatuses. Applications of this detector will be discussed in the last section of this paper. Since event generators like GENIE with anti-neutrino and neutrino interactions are not designed for reactor or solar sources, we were forced to use a power-law and a pair of piecewise continuous exponentials as our two estimating functions, which gave us the ability to extend the cross-section below the calculated data down to the interaction threshold. This was done by using GENIE to calculate the cross sections of 320 different isotopes using GENIE’s entire suite of available physics models.
For each of these isotopes, we next checked the energy threshold for the transition and searched for low-energy excited states which could be used to produce a delayed-coincidence gamma ray for a double-pulse signal. At the same time, the data was extrapolated down to each isotope’s threshold energy using the aforementioned power-law and piecewise exponential functions to determine the cross-sections [7]. After applying these tests to the isotopes in Table 2, the remaining high cross section isotopes were chosen to study further. These isotopes consist of Ba, Gd, and W. Though anti-neutrino detection can occur with all three isotopes, W produced the highest cross section when extrapolated down to reactor anti-neutrino energies. These extrapolated results are shown in Figure 3. As the graphs indicate, the cross-sectional data for W is very similar for both the power and piece-wise exponential functions. Along with having a relatively high cross-section at lower energies and a lower threshold than Ba or Gd, W also has a unique property that does not require the standard neutron capture to create the double pulse coincidence as will be discussed in section 3.
3 Use of a Tungsten isotope as a unique anti-neutrino detector media
Typical anti-neutrino detectors are buried deep underground. Daya Bay, for example, had a detector buried under 250 meters of water equivalent [8]. This is done because it can be difficult to discriminate accidental background detections from real signals. The accidentals in anti-neutrino detectors are generally smaller than those in neutrino detectors due to the characteristic prompt positron signal followed by the s delayed neutron signal. This double-pulse signal is what has allowed many anti-neutrino IBD detectors to be built with less passive shielding than their neutrino counterparts. In this section, we will discuss why Tungsten is such an interesting candidate, and how the use of Tungsten could allow for a much tighter timing window, which would improve the background rejection in turn. For this study of anti-neutrinos, the process depicted in Figure 4 is the only process with a high anti-neutrino cross section we were able to find that has a delayed coincidence fast pulse with W. Its parameters are listed in Table 3.
@@ 3 Use of a Tungsten isotope as a unique anti-neutrino detector media
Anti-neutrino detection would require a scintillator with a large amount of Tungsten, and W is only 14% of naturally occurring Tungsten. Looking for possible scintillators, not many use W. The few which do tend to be low light yield and are instead used for Cherenkov detection methods, like PbWO4. These detection methods do not show the double pulse indicator that is required for isolating a clean anti-neutrino signal. This is changing due to recent developments with glass scintillators, in particular Sodium Phosphate Gallium Tungsten (NaPGaW) has emerged with the necessary material properties needed to perform a double pulse detection. This glass also has the advantage that it can be doped with a 10% mol concentrate of W[4].
@@
By using this scintillator, we are able to control the specific Tungsten isotope, which in turn gives us the ability to use this new detection method of W to Ta. This means that the standard use of neutron capture is not needed because the excited state gamma can serve as the pulse following the positron signal. With the 106 half-life, the double pulse coincidence happens quickly enough that the window for background noise is decreased significantly. Such reductions in background noise can be found in Gallium neutrino detectors using an interaction with a very similar half-life [10]. The Tungsten excited-state reaction does have an upper energy range of 4.5 MeV, due to how weakly bounded W is. If W absorbs more than 4.5 MeV of energy, it can cause spontaneous fission. This process is a type of radioactive decay that occurs predominantly in extremely heavy atomic nuclei. In this event, a single nucleus divides into two or more smaller nuclei, accompanied by the emission of several neutrons and a significant release of energy. Whether or not the isotope is going to be within the spontaneous fission range depends on how much energy each particle takes after the interaction. In either case, these signals will be large and distinct for measurement purposes. Further studies need to be performed to determine the nuclear form factors and the amount of energy in this reaction [7].
@@
Any of the proposed detectors using a Tungsten-based scintillator could be improved by adding a thin film of Gd, perhaps using it as the reflector on the scintillator. This would allow a Tungsten-based detector to still access the IBD signal. In total, there are now methods of detection with W. Such a detector could be designed to look for the lower-rate neutron capture signal, the higher rate W double pulse signal, and the instantaneous spontaneous fission signal. A hybrid detector with smaller segmented Gadolinium Aluminum Gallium Garnet (GAGG) crystals that have W glued between the crystals in thin layers is also an acceptable solution which allows access to the neutron capture signal.
This interaction was used in the main focus of the research, which was to determine if it could be possible to track and triangulate an anti-neutrino source position. To accomplish this, GENIE calculations and subsequent extrapolations were used to determine the full detector capabilities. The results as shown in Figures 5-6 below, show that the range of the W detector is significantly greater than standard IBD method, and this range is sufficiently large that any neutrons also produced by the reactor will be stopped before entering the detector. The capability is also highly dependent on the detector size.
Both the best and the worst case were found using the power output of a 400 MW reactor, representative of the typical power output by a nuclear submarine. For the best (worst) case, we used the higher (lower) cross section of the extrapolated results from GENIE, discussed in Figure 3. The large detector used a mass of 350 kTon, on the order of what a fully loaded oil tanker can hold. The small detector used a mass of 1 kTon, on the order of the fiducial volume of the WATCHMAN experiment.
@@
Simulations of a detector made from GAGG surrounded by tungsten shows that we are capable of resolving inverse beta decay double pulse signals, see Figure 7. These simulations, in combination with the calculations above, give us confidence that a Tungsten-based detector is a promising new technology. To push this technology, we are planing to conduct laboratory studies and experiments to constrain the uncertain aspects of a W detector.
@@ 4 Applications and Conclusion
The initial focus of this research was to determine if the tracking of reactors, specifically within submarines, from the anti-neutrinos emitted was possible, and to create a better method of detection for defense applications. By using the cross-sectional data simulated in GENIE, other simulations were conducted to determine both the energy deposited within the detector, and the ranging capabilities of such a detector. The results indicate that by using W as the target for the anti-neutrino, it is possible to detect a source from a greater distance than by any standard detection method [7]. While this was the starting point, further research of this isotope has resulted in a wide variety of applications.
A neutrino detector made with an isotope that produces a double delayed coincidence signal, given both of these signals are clean with very little background and occur in a short elapsed time, could be applied as a highly-sensitive and low background method of identifying anti-neutrino interactions. Our technique using W for anti-neutrinos can be used to improve science with reactor events [7]. This anti-neutrino detection technique could also be used to improve Geo-neutrino detection, which has applications from geological structure, oil, natural gas and mineral searches and even imaging geological structures on the moon, Venus, asteroids or other earth-like planets.

### v3

@@ Abstract
Low energy anti-neutrinos detected from reactors or other sources have typically used the conversion of an anti-neutrino on Hydrogen, producing a positron and a free neutron. This neutron is subsequently captured on a secondary element with a large neutron capture cross-section such as gadolinium or cadmium. With most neutron captures on gadolinium, it is possible to get two or three delayed gamma signals of known energy to occur. Modern experiments can make measurements with timing on the order of 25 ns. Fast electronics like these allow for the possibility of accessing the very fast signals from the nuclear de-excitation of a heavy nucleus following the prompt positron signal, rather than relying on traditional IBD techniques. We have found an isotope of tungsten, 183W that produces tantalum in the ground state at 2.094 MeV or the first excited state at 2.167 MeV. The excited state of 183Ta∗ emits a signature secondary gamma pulse of 73 keV with a 106 ns half-life. This offers a new delayed coincidence technique that can be used to identify anti-neutrinos with lower background noise. This allows for less shielding than required for modern inverse beta decay detectors.
@@ 1 Introduction
Historically the first direct detection of the neutrino was using anti-neutrinos from a nuclear reactor by Cowan and Reines in the 1950s [1]. Nuclear reactors produce many different energies of anti-neutrinos, not just from the fission process itself, but also from the daughter nuclei produced in the fission, see Figure 1 [2]. The initial focus of this research was to determine if the monitoring of the anti-neutrino rate from nuclear reactors, specifically within submarines, from the anti-neutrinos emitted was possible, and to create a better method of detection for defense applications. Our group has recently been working under NASA grants to design a solar neutrino detector. That work uses a double pulse signal from a gallium interaction, which was seminal in the studies of this paper. To improve detection of reactor anti-neutrinos, rather than only using the double coincidence approach used in the detection of solar neutrinos, we have studied elements which could produce 2 or more coincidence signals with anti-neutrinos, which led us to tungsten 183.
The standard process of anti-neutrino detection[1], depicted in Figure 2, is a weak-nuclear force process of an anti-neutrino interacting with a proton:
where the positron annihilates on an atomic shell electron. This makes a distinct signal from the positron annihilation, but the neutron is fast and must slow down to be captured. The neutron’s passage through matter slows it to capturable energies, but it takes 200 s [3]. When the neutron is absorbed on a heavy element such as gadolinium or cadmium, the resultant excited nucleus emits many different possible cases of secondary gammas when it decays, see Table 1 for gadolinium. In the case of gadolinium one such process only involves the emission of a single gamma and the other 11 have two or more gammas emitted with delayed timings between them. This double coincidence was measured in the 1950s with the slow old electronics of that era ( s [1]), but the large >200 s signal delay made that measurement possible. Today with fast electronics it is possible to consider using a double or triple delayed coincidence with smart electronics selection to reject the out of time background signals. These electronics can be combined with fast scintillators. In contrast with the interaction on a lone proton, if the anti-neutrino has enough energy this reaction can occur on a proton in a heavy element.
A key idea studied here was a comprehensive survey of the known atomic isotopes to find the most promising candidates. We calculated anti-neutrino energy threshold and secondary emission parameters for a particularly promising process, which is described in section 2. We conclude in section 3 with a discussion of possible applications for this technology, including scientific, economic, and national security.
Primary | Secondary | Tertiary | Intensity [% 10-2] |
@@
2 Use of a tungsten isotope as a unique anti-neutrino detector medium
When compared with neutrino detectors, the accidentals in anti-neutrino detectors are generally small due to the characteristic prompt positron signal followed by the s delayed neutron signal. This double-pulse signal allows many anti-neutrino inverse beta decay (IBD) detectors to be built with less passive shielding than their neutrino counterparts. In this section, we will discuss why tungsten is such an interesting candidate, and how the use of tungsten could allow for a much tighter timing window, which would improve the background rejection and increase sensitivity in turn.
@@ 2 Use of a tungsten isotope as a unique anti-neutrino detector medium
Anti-neutrino detection would require a scintillator with a large amount of tungsten, and 183W is only 14% of naturally occurring tungsten. Looking for possible scintillators, we find that not many use tungsten. The few which do tend to be low light yield, like PbWO4, are used to detect Cerenkov light and are not sensitive to the neutron signal required for the double pulse signature of IBD. This is changing due to recent developments with glass scintillators. In particular tungsten gallium-phosphate glasses, such as NaPGaW, have emerged with the necessary material properties needed to perform a double pulse detection[6]. These glasses do not contain a neutron capture target, but they could be used to look for the tantalum de-excitation gamma. These glasses also have the advantage that they can be doped with a 10% mol concentrate of 183W. That work created samples in several concentrations, but it is unclear if 10% was the limit or where they chose to stop[6].
@@
By using the NaPGaW glass scintillator, we are able to control the specific tungsten isotope, which in turn gives us the ability to use this new detection method of 183W to 183m1Ta∗ as shown in Figure 3. This means we can use the excited state gamma as the second pulse after the positron annihilation, rather than the neutron capture. With the 106 half-life, the double pulse coincidence happens quickly enough that the window for background noise is decreased significantly. Such reductions in background noise can be found in gallium neutrino detectors using an interaction with a very similar half-life [7]. The tungsten excited-state reaction does have an upper energy range of 4.5 MeV, due to how weakly bounded 183W is. If 183W absorbs more than 4.5 MeV of energy, that energy can induce fission in the tungsten nucleus. This process is a type of induced fission that occurs predominantly in extremely heavy atomic nuclei. In this event, a single nucleus divides into two or more smaller nuclei, accompanied by the emission of several neutrons and a significant release of energy. Whether or not the isotope is going to be within the induced fission range depends on how much energy the tungsten nucleus takes after the interaction. This induced fission will create a signal very different from the nuclear de-excitation. Because of the energy and heavy nuclei involved, it should be possible to use this signal to also detect anti-neutrinos in the same detector with different signal cuts. Further studies need to be performed to determine the cross-section and nuclear form factors of tungsten as well as the amount of energy in the induced fission reaction [8].
@@
Any of the proposed detectors using a tungsten-based scintillator could be improved by adding a thin film of Gd, perhaps using it as the reflector on the scintillator. This would allow a tungsten-based detector to still access the IBD signal. In total, there are now three methods of detection with 183W. Such a detector could be designed to look for the lower-rate neutron capture signal, the higher rate 183W double pulse signal, and the instantaneous induced fission signal. A hybrid detector with smaller segmented Gadolinium Aluminum Gallium Garnet (GAGG) crystals that have 183W glued between the crystals in thin layers is also an acceptable solution which allows access to the neutron capture signal.
@@
Simulations of a detector made from GAGG surrounded by tungsten shows that a fast scintillator like GAGG is capable of resolving inverse beta decay double pulse signals, see Figure 4[9]. These simulations, in combination with the calculations above, give us confidence that a tungsten-based detector is a promising new technology. To push this technology, we are planning to conduct laboratory studies and experiments to constrain the uncertain aspects of a 183W detector.
@@ 3 Applications and Conclusion
As stated in Section 1, we started this research wanting to improve anti-neutrino detection methods for purposes such as national defense and locating nuclear submarines. The results indicate that 183W is a good candidate for a lower noise detector of anti-neutrinos, and a detector with a sufficiently high cross-section could possibly detect nuclear reactors at large distances [8]. While this was the starting point, further research of this isotope has resulted in a wide variety of applications.
An anti-neutrino detector using an isotope that produces a short-lived excited state could be highly-sensitive and low-background. Our technique using 183W for anti-neutrinos can be used to improve science with reactor events [8]. This anti-neutrino detection technique could also be used to improve Geo-neutrino detection. This has applications in natural resource searches and probing geological structure, which is of interest both on earth and on nearby bodies like the moon, Venus, asteroids, and other earth-like planets.